OpenStudy (anonymous):

find domain and range , f(x) = x^2 + 2 , x element of R.

OpenStudy (anonymous):

so.....Domain is given in the question....Can u guess

OpenStudy (anonymous):

infinity ?

OpenStudy (anonymous):

nop... Domain =R

OpenStudy (anonymous):

teach me please !

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

x element of R. it is given

OpenStudy (anonymous):

i knw this is little difficult topic... in chapter 2

OpenStudy (anonymous):

@ashna domain is to find x such that for all values of x the function is defined..

OpenStudy (anonymous):

yeah ..

OpenStudy (anonymous):

Here can you tell the value of x such that the function is not defined??

OpenStudy (anonymous):

becoz there is no rational function or sqrt function

OpenStudy (anonymous):

more over it is given in the question itself

OpenStudy (anonymous):

no i dont understand @waterineyes .. i know explaining well .. but i dont understand .

OpenStudy (anonymous):

For every value of x you will get f(x).. Right @ashna ?? Or you can tell the value of x such that you will get f(x) undefined ??

OpenStudy (anonymous):

no !

OpenStudy (anonymous):

You have to choose values of x from Real numbers.. For every real number you will get f(x) or not ??

OpenStudy (anonymous):

yes we'll get f(x)

OpenStudy (anonymous):

So can you tell me what will be the domain ??

OpenStudy (anonymous):

all real no's !

OpenStudy (anonymous):

Correct..

OpenStudy (anonymous):

yea .. understood :)

OpenStudy (anonymous):

what about range ?

OpenStudy (anonymous):

We represent like this: \[Domain : \mathbb{R}\]

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

For range you will do like: Let f(x) = y..

OpenStudy (anonymous):

can we also write domain as -infinity and +infinity ?

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

So: \[y = x^2 + 2\] So here can you find x in terms of y ??

OpenStudy (anonymous):

Yes you can write that..

OpenStudy (anonymous):

there is another way

OpenStudy (anonymous):

just plug values for x and see wat u get

OpenStudy (anonymous):

yea .. 1^2 + 2 = 3 2^2 + 2 = 6

OpenStudy (anonymous):

For range we have to put for y not for x..

OpenStudy (anonymous):

nop @waterineyes

OpenStudy (anonymous):

Not this way @ashna

OpenStudy (anonymous):

x^2 + 2

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

wat is the min value of this eq

OpenStudy (anonymous):

I said to find y. Subtract 2 both the sides..

OpenStudy (anonymous):

- infinity

OpenStudy (anonymous):

u follow @waterineyes

OpenStudy (anonymous):

If you think you can explain her then go for it @Yahoo!

OpenStudy (anonymous):

why subtract two ?

OpenStudy (anonymous):

if u r not getting that i will tell another easy...way we can just figure out this by seeing

OpenStudy (anonymous):

You have to find x there..

OpenStudy (anonymous):

nop @waterineyes u go on!

OpenStudy (anonymous):

sumone decide !

OpenStudy (anonymous):

lol. follow waterineyes

OpenStudy (anonymous):

See the equation is: \[y = x^2 + 2\] Here you have to find x.. How can you find x here??

OpenStudy (anonymous):

Don't find the values of x just find x in terms of y..

OpenStudy (anonymous):

by substituting any real no in x ?

OpenStudy (anonymous):

nuuuuu

OpenStudy (anonymous):

No no.. Just in terms of y..

OpenStudy (anonymous):

Subtract 2 both the sides..

OpenStudy (anonymous):

\[x = \sqrt{y-2}\]

OpenStudy (anonymous):

^ trigger happy, mate?

OpenStudy (anonymous):

\[y = x^2 + 2\] Subtract 2: \[y - 2 = x^2\] Take square root: \[x = \sqrt{y-2}\]

OpenStudy (anonymous):

okay .. understood and thanks all !

OpenStudy (anonymous):

so range is (2, infinity ) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now one thing you must remember that the value inside the square root bracket must not be negative: So: \[y - 2 \ge 0\] yes.. Well Done..

OpenStudy (anonymous):

but [2 , infinity

OpenStudy (anonymous):

dont forgot which bracket to use

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

[2 , infinity)

OpenStudy (anonymous):

Note the brackets @ashna \[Range : [2, \infty)\]

OpenStudy (anonymous):

yeah okay :)

OpenStudy (anonymous):

Because this will include 2 also..

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

how about the set notation?

OpenStudy (anonymous):

Meaning ??

OpenStudy (anonymous):

the set-builder notation

OpenStudy (anonymous):

Do you know about that???

OpenStudy (anonymous):

we used to write this in roster form @panlac01

OpenStudy (anonymous):

easy to mention in roster than in set builder

OpenStudy (anonymous):

\[\{x: - \infty < x < \infty\}\] Like this ??

OpenStudy (anonymous):

yes, something like that \[\left\{ x|x \epsilon \mathbb{N} , x<10\right\}\]

OpenStudy (anonymous):

in that case it is >=2

OpenStudy (anonymous):

or I just need to sleep...